OpenStudy (caitlinnr14):

Solve for x: 3(x + 1) = −2(x − 1) + 6

OpenStudy (caitlinnr14):

very confused

OpenStudy (anonymous):

1/5

OpenStudy (caitlinnr14):

do you use distributive property?

OpenStudy (anonymous):

First use distribution 3(x +1) = -2(x - 1) + 6 3 times x 3 times 1 3x + 3 = -2(x - 1) + 6 -2 times x - 2 times -1 3x + 3 = -2x + 2 + 6

OpenStudy (anonymous):

Can you get it from 3x + 3 = -2x + 2 + 6

OpenStudy (caitlinnr14):

answer is 1?

OpenStudy (anonymous):

correct

OpenStudy (caitlinnr14):

thank you

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

Here is a good link to check your work and it helps build confidents in your math http://www.wolframalpha.com/input/?i=3x+%2B+3+%3D+-2x+%2B+2+%2B+6